The equilibrium constant for the reaction N2O4 ⇌ 2 NO2 is 8.75 x 10-2. At equilibrium, a vessel contains N2O4 at a concentration of 1.72 x 10-2 M. What is the concentration of NO2?
1
Expert's answer
2020-05-08T14:10:37-0400

Keq = C(NO)2/C(N2O4),


So:

C(NO) = "\sqrt{\smash[b]{C(N_2O_4)*K_{eq}}}" = "\sqrt{\smash[b]{8.75*10^{-2}*1.72*10^{-2}}}" = 3.88 * 10-2 M


Answer: concentration of NO2 is 3.88 * 10-2 M.